OpenStudy (anonymous):

A man flies a small airplane from Fargo to Bismarck, North Dakota --- a distance of 180 miles. Because he is flying into a head wind, the trip takes him 2 hours. On the way back, the wind is still blowing at the same speed, so the return trip takes only 1 hour 12 minutes. What is his speed in still air, and how fast is the wind blowing?

OpenStudy (anonymous):

Change 1 hr 12 min: 1 + 12/60 = 1.2 hrs : Let s = his speed in still air Let w = speed of the wind then (s-w) = effective speed against the wind and (s+w) = effective speed with the wind : Write a distance equation for each way; dist = time * speed 2(s - w) = 180 1.2(s+w) = 180 Simplify both equations, divide the 1st eq by 2, the 2nd eq by 1.2; results: s - w = 90 s + w = 150 addition eliminates w, find s 2s = 240 s = 240/2 s = 120 mph is the speed in still air Find w s + w = 150 120 + w = 150 w = 150 - 120 w = 30 mph is the speed of the wind Check in 1st original equation 2(120-30) = 180 2(90) = 180

OpenStudy (anonymous):

r, t and d is rate, time and distance respectively. \[\left\{r*t=d, t=\frac{d}{r}\right\} \]\[\text{Solve}\left[\left\{2=\frac{180}{r-w},1+\frac{12}{60}=\frac{180}{r+w}\right\}\right] \]\[\{r\to 120, w\to 30\} \]